OpenStudy (anonymous):

simplify 9^1/2 * 9^1/2

OpenStudy (anonymous):

@PeterPan is it 81 because the fractions equal 1 so its just simple multiplying.

OpenStudy (anonymous):

This is tricky... But remember the laws of exponents... when you multiply two things with the same bases, just add their exponents... \[\Large a^ma^n = a^{m+n}\] So... \[\huge 9^{\frac12}\cdot 9^{\frac12}=9^{\frac12+\frac12}\] And we know two halves make a whole, so...\[\huge 9^{\frac12}\cdot 9^{\frac12}=9^{\frac12+\frac12}=9^1\]

OpenStudy (anonymous):

Thats crazy talk peterpan. but i trust you.

OpenStudy (anonymous):

so its 9

OpenStudy (anonymous):

dont leave meee

OpenStudy (anonymous):

im stranded.

OpenStudy (anonymous):

Sorry... yes, it's 9 :)

OpenStudy (anonymous):

okay thank you. i just didnt want to be wrong. and fail.

OpenStudy (anonymous):

I mean, when you really think about it, \[\Large 9^{\frac12}= \sqrt 9 = 3\] Effectively... \[\Large 9^{\frac12}\cdot 9^{\frac12}=3\cdot 3= 9\]

OpenStudy (anonymous):

By the way, that MIGHT come in handy for you, to know by instinct that square root is exactly the same as raising to an exponent 1/2

OpenStudy (anonymous):

so if its 3 the its 3.5

OpenStudy (anonymous):

And equivalently, cube root is raising to an exponent 1/3, fourth root is raising to an exponent 1/4.... etc

OpenStudy (anonymous):

I'm sorry, what?

OpenStudy (anonymous):

Where did the 5 come from?

OpenStudy (anonymous):

three and a half.

OpenStudy (anonymous):

yeah, why? Why 3 and a half?

OpenStudy (anonymous):

What do you mean by "if its 3..." ?

OpenStudy (anonymous):

the exponient is 3

OpenStudy (anonymous):

Oh, there is no 3 exponent... \[\Large base^{exponent}\]

OpenStudy (anonymous):

so raising it a half would be 3.5 correct?

OpenStudy (anonymous):

logically.

OpenStudy (anonymous):

You mean... \[\Large 3^{\frac12 }=3.5\]?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

Then no. :P \[\Large 3+\frac12 = 3.5\qquad\color{green}{\checkmark}\] \[\Large 3^{\frac12 }= 3.5 \qquad \color{red}{\bf{\times}}\]

OpenStudy (anonymous):

\[\Large 3^{\frac12}= \sqrt3 \approx 1.732\]

OpenStudy (anonymous):

the demnoiter 2 so it would be 3*3

OpenStudy (anonymous):

No... \[\Large 3\cdot 3 = 3^2 \ne 3^{\frac12}\]

OpenStudy (anonymous):

In other words, 3*3 is 3^2 and NOT 3^(1/2)

OpenStudy (anonymous):

so if its 3 1/2 i divide 3 by 2

OpenStudy (anonymous):

Multiplying the exponent to the base? \[\Large 3^{\frac 12}= 3\times \frac12 = \frac32 \qquad \color{green}{?}\]

OpenStudy (anonymous):

My teacher calls that step the "never ever do that or I'll kill you" step... you don't multiply the exponent to the base...

OpenStudy (anonymous):

Just take it from me, the exponent 1/2 means you take the square root of the base... \[\huge a^{\frac12}= \sqrt a\]

OpenStudy (anonymous):

and nothing else...

OpenStudy (anonymous):

so the 1 just vanishes?

OpenStudy (anonymous):

Well, technically \[\huge a^{\frac12}= \sqrt{a^1}\] but, clearly, a^1 is just a.

OpenStudy (anonymous):

okay what if (8x^-6) ^1/3

OpenStudy (anonymous):

Okay, two new rules come into play, first, when you raise the product of two things to an exponent, you just distribute the exponent over the product... \[\Large (ab)^m = a^mb^m\]

OpenStudy (anonymous):

So, with that in mind, we get \[\huge (8x^{-6})^{\frac13 }= 8^{\frac13}\cdot (x^{-6})^{\frac13}\] yeah?

OpenStudy (anonymous):

so 8 cubed times -6x cubed

OpenStudy (anonymous):

nu-uh.... does that look like a cube to you? THIS  is what a cube looks like :D no, seriously, a cube is an exponent of 3, NOT 1/3 an exponent of 1/3 means cube ROOT.

OpenStudy (anonymous):

So, what is the cube root of 8? :)

OpenStudy (anonymous):

2.66666 or 2.7

OpenStudy (anonymous):

<whistles> that's the square root (read in an annoying childish sing-song voice)

OpenStudy (anonymous):

face palm

OpenStudy (anonymous):

So... what IS the cube root of 8, then? :)

OpenStudy (anonymous):

8 divided by 3 right???

OpenStudy (anonymous):

you're doing it again... multiplying the exponent to the base.... my teacher's gonna kill you :D

OpenStudy (anonymous):

shh,. dont tell on me

OpenStudy (anonymous):

lol, in all seriousness, the exponent isn't something you multiply (or divide) with the base... they're off limits from each other, if you know what I mean...

OpenStudy (anonymous):

The cube root of 8, is that number, which, when you multiply to itself three times, you get 8. r x r x r = 8

OpenStudy (anonymous):

512

OpenStudy (anonymous):

dont tell me im wrong again...

OpenStudy (anonymous):

Well, you're not right :P You really need to see that distinction between cube and cube ROOT. LOL The cube of 8 is 512, because 8 x 8 x 8 = 512... however, the cube ROOT of 8 is that number such that when you multiply to itself three times, you get 8. Take for example, 27, like earlier... the cube ROOT of 27 is 3, because 3x3x3 = 27 ^.^

OpenStudy (anonymous):

so its 24.

OpenStudy (anonymous):

And you tried dividing 8 by the exponent, 1/3... -.-

OpenStudy (anonymous):

No... suffice it to say, the cube root of 8 is 2... because \[\Large 2^3 = 2\times 2\times2 = 8\]

OpenStudy (anonymous):

no. draw it out for me peterpan.

OpenStudy (anonymous):

So, everything understood, the cube root of 8 is 2?

OpenStudy (anonymous):

2.66666

OpenStudy (anonymous):

You multiplied 8 by 1/3 again?

OpenStudy (anonymous):

Pull your act together :D If in doubt, check if 2.66666 is indeed the cube root of 8, by *cubing* it... 2.66666^3 is like, 18.96.... definitely not 8 :P

OpenStudy (anonymous):

OOOOOOOOHHHHHHH!!!!!! that waht i wasnt putting together.

OpenStudy (anonymous):

so... the cube root of 8?

OpenStudy (anonymous):

eight divided by 3

OpenStudy (anonymous):

^3

OpenStudy (anonymous):

You multiplied 8 by 1/3 again... three strikes, you're out :D

OpenStudy (anonymous):

i got 72

OpenStudy (anonymous):

no i divided.

OpenStudy (anonymous):

i got 2.66666

OpenStudy (anonymous):

dividing by 3 is the same as muliplying by 1/3. and either way, it's wrong... come on, Chris, how many times do I have to tell you that you don't multiply or divide the base with the exponent?

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

\[\Large 8^{\frac13}\ne 8\div 3\]\[\large \color{red}{!!!!!!!!!!!}\]

OpenStudy (anonymous):

so its 8x8x8?

OpenStudy (anonymous):

<awkward smile> no.... \[\Large 8\times 8\times 8 = 8^3 \ \ne \ 8^{\frac13}\]

OpenStudy (anonymous):

so what is 8 1/3

OpenStudy (anonymous):

I actually already dropped it :) \[\Large 8^{\frac13}= 2\] because.... \[\huge 2\times2\times2 = 8\]

OpenStudy (anonymous):

so its backwards sorta say.

OpenStudy (anonymous):

so its 2/x^2

OpenStudy (anonymous):

hmmmn....I realise that doing things backwards is harder than forwards... (like how I struggled with subtraction after mastering addition ^.^) But it's something you need to learn :P And unfortunately, this is only part of it... \[\huge (8x^{-6})^{\frac13 }= \color{red}{8^{\frac13}}\cdot (x^{-6})^{\frac13}=\color{red}2\cdot (x^{-6})^{\frac13}\]

OpenStudy (anonymous):

Now we need to deal with this bit... \[\huge 2\cdot\color{blue} {(x^{-6})^{\frac13}}\]

OpenStudy (anonymous):

Which brings us to another law of exponents... \[\huge (a^m)^n= a^{mn}\] When you raise a power to another power, you multiply the exponents... So how do you handle this \[\huge 2\cdot\color{blue} {(x^{-6})^{\frac13}}\]

OpenStudy (anonymous):

i would multiply -6 * 1/3

OpenStudy (anonymous):

That's right... notice the difference... you may multiply exponents with other exponents, but NEVER with the base...

OpenStudy (anonymous):

so -6/3?

OpenStudy (anonymous):

mhmm... simplify?

OpenStudy (anonymous):

-2?

OpenStudy (anonymous):

BINGO . \[\huge 2\color{blue}{x^{-2}}\]

OpenStudy (anonymous):

So, is that in your choices?